i am doing solving systems by elimination and i don't understand how to the problem. it is 3x-y=-2 and -2x+2y=8
Answer by richwmiller(17219)

(1) 3x-y=-2
(2) -2x+2y=8
multiply 3x-y=-2 by 2
(1a) 6x-2y=-4
add (1a) to second (2) equation
-2x+2y=8
6x-2y=-4
4x=4
x=1
substitute x=1 into any of the three equations
-2x+2y=8
-2+2y=8
add 2
2y=10
y=5
solution
x=1
y=5